Area of Surface of Revolution

1. 9𝑦 = π‘₯^2 , 0 ≀ π‘₯ ≀ 2 ; x-axis

2. 𝑦 = 𝑙𝑛(π‘₯^2 βˆ’ 1) , 2 ≀ π‘₯ ≀ 3 ; y-axis

3. 𝑦 = 3^√π‘₯ , 0 ≀ π‘₯ ≀ 1 ; y-axis

4. 𝑦 = √9 βˆ’ π‘₯^2 , βˆ’2 ≀ π‘₯ ≀ 2 ; x-axis


1
Expert's answer
2021-05-25T18:18:13-0400

1. 9𝑦 = π‘₯2 , 0 ≀ π‘₯ ≀ 2 ; x-axis ..............Equation(1)


Area of the surface is given by


A = "\int" 02 2 "*\pi" "*" y "\sqrt{1 + (\dfrac{dy}{dx})^2}" dx



Differentiating Equation(1) with respect to x, we have


"\dfrac{dy}{dx}" = 2x / 9


So, area is


A = "\int" 02 2"*\pi" "*" "\dfrac{x^2}{9}" "\sqrt{1 + \dfrac{4x^2}{81}}" dx ........................Equation(2)



A = "\dfrac{4}{81}" "\pi" "\int" 02 x2"\sqrt{x^2 + \dfrac{81}{4}}" dx


Let x = "\dfrac{9}{2}\tan\theta"


dx = (9/2) sec2 "\theta" d"\theta"


A = "\dfrac{4}{81}" "\pi" "\int" 0 "\phi" "\dfrac{81}{4} tan^2\theta" "\dfrac{9}{2}\sec\theta" (9/2) sec2 "\theta" d"\theta" ..............[ "\phi" = arctan(4/9)]


A = "\pi" "\int" 0 "\phi" "\ tan^2\theta" "\dfrac{9}{2}\sec\theta" "\dfrac{9}{2}" sec2 "\theta" d"\theta"


A = "\dfrac{81\pi}{4}" "\int"0 "\phi" "tan^2\theta sec^3\theta" d"\theta"


On solving the above integral we get


A = 2.43 square units.




2. 𝑦 = 𝑙𝑛(π‘₯^2 βˆ’ 1) , 2 ≀ π‘₯ ≀ 3 ; y-axis ...............Equation(2)


x varies from 2 to 3

So, y varies from ln 3 to ln 8


Area of the surface is given by,


A = "\int" ln 3 ln 8 2 "*\pi" "*" x "\sqrt{1 + (\dfrac{dx}{dy})^2}" dy


Differentiating Equation (2) with respect to y we have


"\dfrac{dx}{dy} = \dfrac{e^y}{2\sqrt{e^y + 1}}"


A = "\int" ln 3 ln 8 2 "*\pi" "*" "\sqrt{e^y + 1}" "*" "\dfrac{e^y + 1}{2\sqrt{e^y + 1}}" dy


A = "\int" ln 3 ln 8 "*\pi" (ey + 1 )dy


A = "\pi" [(ey + y ] ln 3 ln8


A = 18.78 square units.




3. 𝑦 = 3^√π‘₯ , 0 ≀ π‘₯ ≀ 1 ; y-axis ................Equation(3)


x = y3

dx/dy = 3y2


The area is given as


A = "\int" 0 1 2 "*\pi" "*" x "\sqrt{1 + (\dfrac{dx}{dy})^2}" dy

A = "\int" 0 1 2 "\pi" y3 "\sqrt{1 + 9y^4}" dy


Let 1 + 9y4 = t

36 y3 dy = dt


A = "\int" 1 10 2 "\pi" "\dfrac{\sqrt{t}}{36}" dt


A = "\dfrac{2\pi}{36} * \dfrac{2}{3} \int" 1 10 [ t1.5 ]1 10


A = 3.561 square units.





4. 𝑦 = √9 βˆ’ π‘₯^2 , βˆ’2 ≀ π‘₯ ≀ 2 ; x-axis ...........................Equation(4)


Area of the surface is given by


A = "\int" -2 2 2 "*\pi" "*" y "\sqrt{1 + (\dfrac{dy}{dx})^2}" dx



Differentiating Equation(1) with respect to x, we have


"\dfrac{dy}{dx}" ​= "\dfrac{- x}{\sqrt{9 -x^2}}"


So, area is



A = "\int" -2 2 2 "*\pi" "*" "\sqrt{9-x^2} * \dfrac{3}{\sqrt{9 - x^2}}" dx


A = "\int" -2 2 6"\pi" dx


A = 24 "\pi" square units


A = 75.36 square units.